OpenStudy (anonymous):

Larry and Peggy are making decisions about their bank accounts. Larry wants to deposit $350 as a principle amount, with an interest of 4% compounded quarterly. Peggy wants to deposit $350 as the principle amount, with an interest of 6% compounded monthly. Explain which method results in more money after 2 years. Show all work.

OpenStudy (anonymous):

I know the formula is A = P(1 + r/n)nt but i'm still a little confused on how to solve this. Help is appreciated =)

OpenStudy (anonymous):

You need to enter the known values into the formula. Which values do you already know?

OpenStudy (anonymous):

The rate of interest, the time and the principle?

OpenStudy (anonymous):

Great, so how does the formula look when you input that information?

OpenStudy (anonymous):

A = 350(1 + 0.04/4)(4 * 2)? A = 350(1 + 0.06/12)(4 * 2)?

OpenStudy (anonymous):

Let's look at the first one. A=350 x (1+ (.04/4))^4 Try entering it like this in the calculator you should get 364.21

OpenStudy (anonymous):

Okay, so that's how much Larry would make?

OpenStudy (anonymous):

I just noticed that this is for two years. A=350 x (1+ (.04/4))^4*2 =?

OpenStudy (anonymous):

379?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And then Peggy's method would result with 479 dollars?

OpenStudy (anonymous):

no, I have a different answer for her

OpenStudy (anonymous):

350 x (1+(.06/12))^(12*2)

OpenStudy (anonymous):

Ohhhhhh, 396 dollars?

OpenStudy (anonymous):

You still seem to be a little off, but closer

OpenStudy (anonymous):

395? Haha

OpenStudy (anonymous):

ok. You can round to 395, but it is 394.51

OpenStudy (anonymous):

Do you understand every thing that is going on? Do you have any more questions on this problem?

OpenStudy (anonymous):

No I think I can word it properly now, thanks! If you don't mind though, I still have another question I need help with

OpenStudy (anonymous):

ok, let's give it a try

OpenStudy (anonymous):

There's some graphs as well so i'll copy and paste and then clip in the pictures

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Kylee manages a small theme park and she has been analyzing the attendance data. Kylee finds that the number of visitors increases exponentially as the temperature increases, and this situation is represented by the function f(x) = 4^x. Kylee also finds a linear equation that models the number of people who leave the park early depending on the change in temperature, and it is represented by f(x) = −x + 5. The graph of the two functions is below. Find the solution to the two functions and explain what the solution represents.

OpenStudy (anonymous):

@satellite73 Maybe you could help me with this?

OpenStudy (anonymous):

hmm, that one is a bit harder. I can explain how the number are located on the graph, but I'm not sure how to solve the problem. This change in rate is a calculus problem, which I have done, but none where the exponent was the variable. 4^x seems off. If it were x^4 you would bring the four out front and subtract one from the exponent 4x^3. Did this some how ask for the fog(x)?

OpenStudy (anonymous):

I'm not exactly sure, I just know that the solution represents the intersection of the 2 functions. It might be a type of system of equations? I copied and pasted the question exactly

OpenStudy (anonymous):

My guess is that for every x degrees five people show up and one person leaves. Thus there is a net of four people

OpenStudy (anonymous):

I don't think that applies to what i've been learning. I'm completely stumped too, somehow I have to word this into an essay response

OpenStudy (anonymous):

Sorry, I don't know.

OpenStudy (anonymous):

Alright, thanks for your help anyways! I'll post this as a separate question and hopefully someone else knows.